OpenStudy (anonymous):

what is the distance from (-7,4) to (-9,2)?

OpenStudy (anonymous):

distance= \[\sqrt{(-7+9)^2+(4-2)^2}\]

OpenStudy (anonymous):

sum of the square of the differnce between x and y coordinates of the points

OpenStudy (anonymous):

do the square and the square root cancel each other out?

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

inside the sqrt there r two terms adding up

OpenStudy (anonymous):

simplify those first

OpenStudy (anonymous):

sqrt(4+4)=sqrt(8) =2sqrt(2)

OpenStudy (shadowfiend):

Right, remember you can't cancel two squares that are added together with a square root; only if they're multiplied together :)

OpenStudy (anonymous):

The answer is \[2\sqrt{2} ?\]